# Scaling LLM Planning: NL2FLOW for Parametric Problem Generation and Rigorous Evaluation

### v1

@@ 3.4 Language Model Response Parsing
To establish a lower bound on large language model performance, I employed minimal parsing techniques for both translation and planning tasks, prioritizing simplicity and avoiding potentially biasing post-processing steps. For the translation task, model outputs were structured as JSON strings representing a planning problem. These strings were then parsed using PyDantic, a data validation and settings management library, to ensure adherence to a predefined schema. This structured format enabled automated evaluation metrics and facilitated downstream analysis. For the planning task, each action within a generated plan was extracted and stored as a string within a Python list. This list was subsequently used as input to the NL2FLOW debugger, a symbolic planning validation tool that assesses the logical consistency and feasibility of the generated plan.
3.5 Plan Validation
@@
To assess the soundness of generated plans, the pddl compilation omits goal achievement. My validation approach distinguishes between soundness and validity. For soundness checking, I compiled the planning problem into PDDL, but instead of including the original goals as requirements, I incorporated the actions reported as executed within the LLM’s response. This approach tests the LLM’s claimed actions actually executable. Validity is checked by compiling with both original goals and actions in LLM’s plan as requirements; failure to achieve goals indicates a non-valid plan where a planner finds a longer optimal plan.
@@
To quantify plan optimality, I compare the length of the LLM-generated plan to the length of optimal plans derived using the $\ast$ planner. I formulate the planning problem in PDDL using only the original goals as requirements and leverage $\ast$ to obtain a set of optimal plans. Following a successful validity check, an LLM-generated plan is considered sub-optimal if its length exceeds the length of the shortest $\ast$-derived optimal plan. This comparison provides a quantitative measure of how efficiently the LLM solves the planning problem relative to a state-of-the-art planner.
I investigated the use of VAL [16] for plan validation; however, VAL does not currently support the PDDL language features necessary to represent the complexities introduced by my NL2FLOW compilation process, precluding its use in this evaluation.
@@ 4.1.1 Plan Length
Figure 3: The distribution of plan lengths (number of steps) generated by LLMs
@@ 4.1.2 Sound, Valid, and Optimal Planning
Overall, Llama-3.3-70B-instruct, when paired with the verbose prompt template, demonstrated the strongest planning performance (Figure 4). This model achieved the highest rates of generating sound plans (80%), valid plans (77%), and optimal plans (64%) among all tested models. Utilizing the concise prompt template resulted in a noticeable decrease in performance, with rates dropping to 71%, 60%, and 46% respectively. A similar, though more pronounced, trend was observed for Llama-3.1-8B-instruct. For clarity, a sound plan is defined as one that is executable; a valid plan is executable and achieves the goal state; and an optimal plan is executable, reaches the goal state, and does so with minimal cost.
@@ 4.1.3 No Plan
A key limitation observed across all models was the difficulty in correctly identifying unsolvable problems (Figure 5) – those for which no plan exists. The superior performance of Llama-3.3-70B-instruct in generating sound, valid, and optimal plans was largely attributable to its enhanced ability to recognize these ‘no plan’ scenarios. Analysis of explanatory variables – including the number of actions, arity, coupling, the proportion of slot-fillable variables, and the number of goals – revealed differing factors contributing to failures in identifying ‘no plan’ problems depending on the prompt template (Tables LABEL:tab:lrnoplanlong and LABEL:tab:lrnoplanshort). With the verbose prompt template, arity was the sole consistent predictor of failure (higher arity correlated with a higher rate of incorrectly attempting to generate a plan when none existed). Conversely, using the concise prompt template, the number of goals was the primary factor, with a higher number of goals increasing the likelihood of failing to identify ‘no plan’ problems.
Figure 4: Comparison of LLM performance in generating sound, valid, and optimal plans using verbose and concise prompt templates
Figure 5: Comparison of large language model performance (accuracy) in detecting ’no plan’ scenarios, evaluated with verbose and concise prompt templates
@@ 4.1.4 Feasible Plan
[…]
Regarding optimal plan generation, coupling consistently explained a significant portion of the variance across all models when the verbose prompt was used (Table LABEL:tab:lrplanoptimallong). This trend was observed only in Llama-3.3-70B-instruct when using the concise prompt (Table LABEL:tab:lrplanoptimalshort). These results suggest that the impact of problem characteristics on plan generation is both model- and prompt-dependent.
Figure 6: Performance comparison of large language models on planning tasks with solutions
@@ 4.2 Natural Language to JSON Translation
Llama-3.1-405B-Instruct-FP8 achieved the highest success rate in generating valid JSON strings (99%) and minimizing errors in content (Figure 7). Specifically, 73% of the JSON outputs from this model contained all required elements without hallucinated information. Codellama-34B-Instruct-HR and Llama-3.3-70B-Instruct consistently failed to generate valid JSON. Across all models, the primary source of translation error was the omission of required elements, rather than the inclusion of hallucinated information. Linear regression analysis revealed a negative correlation between the proportion of correctly translated elements and both the number of actions and their arity (Table LABEL:tab:olstranslation). Conversely, a positive correlation was observed between correct translation and both the degree of coupling between actions and the proportion of slot-fillable variables.

### v2

@@ 2.4 Language Model Response Parsing
To establish a lower bound on large language model performance, I used minimal parsing techniques for both translation and planning tasks, prioritizing simplicity and avoiding potentially biasing post-processing steps.
For the translation task, model outputs were structured as JSON strings representing a planning problem. These strings were then parsed using PyDantic, a data validation and settings management library, to ensure adherence to a predefined schema. This structured format enabled automated evaluation metrics and facilitated downstream analysis. The initial intermediate representation schema used nested structures to represent custom actions; however, this complexity proved unnecessary, as the same information could be represented with a simplified schema. To evaluate the capacity of LLMs to translate natural language descriptions of workflow problems into JSON, I used a streamlined schema (Listing 9) defined as follows:
  • $actions$ represents custom actions.
  • $mappings$ corresponds to the ’mappings’ present in the intermediate representation.
  • $available\_data$ is a list of variables with known values at the workflow’s initiation.
  • $askable\_parameters$ is a list of slot-fillable variables for user input.
  • $unaskable\_parameters$ is a list of variables for which the system is prohibited from prompting the user.
  • $goal\_action\_ids$ corresponds to the ’goal_action_ids’ in the intermediate representation.
Listing 9: JSON schema used for evaluating LLM’s translation capacity
actions”: [
id”: a__1”,
input”: [
v__1”,
v__2
output”: [
v__3
mappings”: [],
available_data”: [
v__5”,
askable_parameters”: [
v__1”,
unaskable_parameters”: [
v__4”,
goal_action_ids”: [
a__3
An example of this schema, used for evaluating LLM translation capacity, is provided as ‘translation_object.json‘ in the supplementary material.
For the planning task, each action within a generated plan was extracted and stored as a string within a Python list. This list was subsequently used as input to the NL2Flow debugger, a symbolic planning validation tool that assesses the logical consistency and feasibility of the generated plan.
2.5 AI Planning
AI planning addresses the problem of finding a sequence of actions that transforms an initial state into a goal state. Formally, a planning problem is defined as $P=\langle P,F,D,I,G\rangle$, where:
  • $P$ is a finite set of predicates describing properties of the world.
  • $F$ is a finite set of numeric variables.
  • $D$ is the domain model, comprising a set of action schemas.
  • $I$ is the initial state, describing the world at the start of planning.
  • $G$ is the goal state, specifying the desired outcome.
I adopt the closed-world assumption, meaning any fact not explicitly stated in the initial state is considered false. An action $a$ is defined by its preconditions and effects: $a=\langle pre(a),eff(a)\rangle$. Preconditions, $pre(a)$, define the conditions that must hold in a state for the action to be applicable. Effects, $eff(a)$, specify how the state changes upon execution of the action.
A state $s$ is a conjunction of predicates and numeric variable assignments: $s=\langle P_{s},V_{s}\rangle$, where $P_{s}\subseteq P$ represents the set of true predicates, and $V_{s}\in\mathbb{Q}^{|F|}$ represents the assignment of rational values from $\mathbb{Q}$ to each numeric variable in $F$.
[…]
An effect $eff(a)$ is a triple $\langle P^{+}_{eff},P^{-}_{eff},U_{eff}\rangle$, where:
  • $P^{+}_{eff}\subseteq P$ is the set of predicates that become true.
  • $P^{-}_{eff}\subseteq P$ is the set of predicates that become false.
  • $U_{eff}$ is a set of numeric updates, each of the form $\langle f,:=/+=/-=,exp\rangle$, where $f\in F$ is the updated variable and $exp$ is the new value or change in value.
Applying an action $a$ in state $s$ is possible if all predicates in $pre(a)$ are present in $P_{s}$ and all numeric constraints in $pre(a)$ hold true when evaluated using $V_{s}$. The resulting state $s^{\prime}=\langle P^{\prime}_{s},V^{\prime}_{s}\rangle$, is computed as follows: $P^{\prime}_{s}=(P_{s}\setminus P^{-}_{eff})\cup P^{+}_{eff}$, and each variable $f\in F$ is updated in $V^{\prime}_{s}$ according to the operator and expression specified in $U_{eff}$.
A plan $\pi$ is a sequence of actions. The trajectory of a plan, starting from initial state $I$, is the sequence of states $s_{0},s_{1},\dots$
2.6 PDDL Compilation
This study used NL2Flow to translate an intermediate representation of a workflow creation problem into PDDL. The PDDL formulation within NL2Flow enabled cost-based planning to identify efficient workflows (plans). While NL2Flow was designed to represent a broad range of workflow automation concepts, this study focused on generating PDDL for simple, low-difficulty workflows. For instance, the mapping capacity feature, which allowed variable values to be shared, was excluded from the current implementation.
Planning problems in this work were formalized using the Planning Domain Definition Language (PDDL). PDDL separated each problem’s description into a domain model, which defined the possible actions and their effects, and a problem instance, specifying the initial state and goal conditions. Example domain and problem files used in this study (domain.pddl and problem.pddl, respectively) were provided as supplementary material. A domain model included declarations of predicates, functions, and actions, while a problem instance defined the initial state through a set of grounded predicates and function assignments. Any fact not explicitly stated as true in the initial state was considered false. Grounding involved replacing variables in predicates and functions with concrete values.
2.6.1 Types
The domain defines several types:
  • $generic$: Represents the objects being processed – the entities whose slots need filling.
  • $operator$: Represents actions the agent can take.
  • $has-done-state$: Represents temporal states (past, present, future) to track whether an action has been performed.
  • $constraint-status$: Represents certainty levels (True, False) likely used in constraints or conditions.
  • $datum-state$: Represents the certainty of information about an object (certain, uncertain, unknown).
  • $label$: (Currently unused) labeling objects or slots.
  • $object$: A base type for extensibility.
2.6.2 Constants
Many constants are defined for each type. The ’generic’ constants ($v\_\_x$) represent specific instances of the objects being processed. The $has{-}done{-}state$ constants track the temporal status of actions. The $constraint{-}status$ and $datum{-}state$ constants define the possible values for these attributes. $operator$ constants ($a\_\_x$) represent specific, pre-defined actions.
2.6.3 Predicates
The predicates define the relationships and properties within the domain:
  • $(has\_done\>?x\>-\>operator\>?x2\>-\>has{-}done{-}state)$: Indicates that an operator has been performed in a specific time state.
  • $(been\_used\>?x\>-\>generic)$: Marks an object as having been processed.
  • $(new\_item\>?x\>-\>generic)$: Indicates an object is newly introduced.
  • $(known\>?x\>-\>generic\>?x2\>-\>datum{-}state)$: Represents the agent’s knowledge about an object’s certainty. This is central to the planning process.
  • $(not\_slotfillable\>?x\>-\>generic)$: Indicates an object cannot have its slots filled (perhaps due to inherent properties).
[…]
  • $(mapped\>?x\>-\>generic)$: Indicates an object has been mapped.
  • $(not\_usable\>?x\>-\>generic)$: Indicates an object cannot be used.
  • $(mapped\_to\>?x\>-\>generic\>?x2\>-generic)$: Indicates an object is mapped to another object.
  • $(free\>?x\>-\>generic)$: Indicates an object is available for processing.
[…]
2.6.4 Functions
Functions are used to represent numeric state variables. They allow you to model quantities that change over time and are used in numeric effects of actions.
  • $(total{-}cost)$: A numeric function representing the cumulative cost of the plan.
  • $(slot\_goodness\>?x\>-\>generic)$: A numeric function that likely evaluates the q̈ualityör usefulness of filling a slot for a given object.
  • $(affinity\>?x\>-\>generic\>?x2\>-\>generic)$: A numeric function that likely measures the similarity or relatedness between two objects.
2.6.5 Actions
The domain defines two action types. Custom actions are named according to the pattern $a\_\_x$.
  • $a\_\_x$:
    • Precondition: Require that the action hasn’t been done in the past and that specific objects ($v\_\_x$) are known with certainty.
    • Effect: Mark the action as done in the present. Mark the object as used. Free up objects. Mark the newly freed object as known with certainty. Mark the newly freed object as not mapped. Increase the $total{-}cost$ by 10.
Listing 10: ’custom’ action example in PDDL
(:action a__2
:parameters ()
:precondition (and (not (has_done a__2 past)) (known v__7 certain))
:effect (and
(has_done a__2 present)
(been_used v__7)
(free v__9)
(known v__9 certain)
(not (mapped v__9))
(increase (total-cost ) 10)
For example, Listing 10 illustrate the structure of ’custom’ actions defined in the domain model. This action, $a\_\_2$, has no parameters. Its precondition requires that the action hasn’t been performed in the past ($not(has\_done\>a\_\_2\>past)$) and that a variable $v\_\_7$ is known with certainty ($known\>v\_\_7\>certain$). The action’s effect marks the action as done in the present ($has\_done\>a\_\_2\>present$), indicates $v\_\_7$ has been used, makes $v\_\_9$ free, establishes $v\_\_9$ as known with certainty, indicates $v\_\_9$ is no longer mapped, and increases the $total\mbox{-}cost$ by 10.
The $ask$ action prompts the system to request information from the user if parameters are slot-fillable (Listing 11).
Listing 11: ’ask’ action example in PDDL
(:action ask
:parameters (?x - generic)
:precondition (and (not (known ?x certain)) (not (not_slotfillable ?x))
:effect (and
(free ?x)
(mapped_to ?x ?x)
(known ?x certain)
(not (not_usable ?x))
(not (mapped ?x))
(increase (total-cost ) (slot_goodness ?x))
  • ‘ask‘:
    • Takes a $generic$ object as a parameter.
    • Precondition: Requires that the object is not currently known with certainty and is slotfillable ($not\>not\_slotfillable$).
    • Effect: Frees the object. Maps the object to itself (a placeholder). Marks the object as known with certainty. Marks the object as usable (not $not\_usable$). Marks the object as not mapped. Increases the $total-cost$ by the $slot\_goodness$ of the object.
2.6.6 Problem
A problem describes a specific instance of a planning problem. It works in conjunction with a domain, which defines the general rules and actions applicable to a class of problems. A problem contains:
  • Initial state: The starting situation for the planner. This is a description of what’s true at the beginning.
  • Goal state: What the planner needs to achieve to solve the problem. This is a condition that must be true when the plan is complete.
  • Specifies objects: The specific objects that exist in this particular problem instance.
In this study, the initial state was defined by a set of variables ($v\_\_x$), each assigned a $slot\_goodness$ value of 1000000 if slot-fillable, and 150000 if not. These differing $slot\_goodness$ values represented a ’soft’ constraint, which was not used in this study because the precondition for the ’ask’ action prevented the system from requesting information for variables that were not slot-fillable. The $mapped\_to$ predicate establishes a self-mapping for each resource, essentially stating that each resource is initially assigned to itself. A ‘new_item‘ predicate indicates the creation of a new resource. The goal is to achieve conditions to complete a task or tasks ($(has\_done\>a\_\_x\>present)and(has\_done\>a\_\_x\>present)$).
2.7 Plan Evaluation
NL2Flow evaluated plans generated by LLMs using the Planning Domain Definition Language (PDDL). LLM-generated plans were extracted from text and converted to PDDL via NL2Flow, using an intermediate representation from the data generator. The resulting PDDL formulations were then assessed with a symbolic planner. For clarity, a sound plan is defined as one that is executable; a valid plan is executable and achieves the goal state; and an optimal plan is executable, reaches the goal state, and does so with minimal cost.
@@ 2.7 Plan Evaluation
To assess the soundness of generated plans, the pddl compilation omits goal achievement. NL2Flow’s approach distinguishes between soundness and validity. For soundness checking, the planning problem was compiled into PDDL, but instead of including the original goals as requirements, the actions reported as executed within the LLM’s response were incorporated. This approach tests the LLM’s claimed actions actually executable. Validity was checked by compiling with both original goals and actions in LLM’s plan as requirements; failure to achieve goals indicates a non-valid plan where a planner finds a longer optimal plan. In this low-difficulty workflow problem setting, a plan was invalid, also, if the system prompted the user for a non-slot-fillable variable, or if a ’custom’ action was invoked with unknown input parameter values. I investigated the use of VAL [5] for plan validation; however, VAL did not support the PDDL language features necessary to represent the complexities introduced by NL2Flow’s PDDL compilation process, precluding its use in this evaluation.
@@
To quantify plan optimality, the length of the LLM-generated plan was compared to that of optimal plans derived using the $\ast$ planner [6]. The planning problem was formulated in PDDL using only the original goals as requirements and use $\ast$ to obtain a set of optimal plans. Following a successful validity check, an LLM-generated plan was considered sub-optimal if its length exceeded the length of the shortest $\ast$-derived optimal plan. This comparison provides a quantitative measure of how efficiently the LLM solves the planning problem relative to a state-of-the-art planner.
Figure 2: Comparison of LLM performance in generating sound, valid, and optimal plans using verbose and concise prompt styles
Following normalization of the explanatory variables (Tables LABEL:tab:lrnoplanlong, LABEL:tab:lrnoplanshort, LABEL:tab:lrplanvaliditylong, LABEL:tab:lrplanvalidityshort, LABEL:tab:lrplanoptimallong, LABEL:tab:lrplanoptimalshort, and LABEL:tab:olstranslation in Appendix), I used logistic regression, implemented in the Python statsmodels package, to assess the contribution of these variables to different aspects of LLM workflow generation capacity. Ordinary least squares (OLS) linear regression, also implemented in statsmodels, was used to estimate the contribution of each variable to the success of natural-language-to-intermediate-representation translation.
@@ 3.1.1 Plan Length
Figure 3: The distribution of plan lengths (number of steps) generated by LLMs with verbose prompt style
Figure 4: The distribution of plan lengths (number of steps) generated by LLMs with concise prompt style
@@
Prompt style did not significantly alter the plan length distribution for most models (Figures 3 and 4). However, the plan length distributions for Llama-3.1-8B-instruct and Mixtral-8x22B-instruct-v0.1 did shift based on prompt style, suggesting these models are more sensitive to prompting variations.
@@ 3.1.2 Sound, Valid, and Optimal Planning
Overall, Llama-3.3-70B-instruct, when paired with the verbose prompt style, demonstrated the strongest planning performance (Figure 2). This model achieved the highest rates of generating sound plans (80%), valid plans (77%), and optimal plans (64%) among all tested models. Using the concise prompt style resulted in a noticeable decrease in performance, with rates dropping to 71%, 60%, and 46% respectively. A similar, though more pronounced, trend was observed for Llama-3.1-8B-instruct.
@@ 3.1.3 No Plan
A key limitation observed across all models was the difficulty in correctly identifying unsolvable problems (Figure 5) – those for which no plan exists. The superior performance of Llama-3.3-70B-instruct in generating sound, valid plans was largely attributable to its ability to recognize these ‘no plan’ scenarios. Analysis of explanatory variables – including the number of custom actions, arity, coupling, the proportion of slot-fillable variables, and the number of goals – showed that factors influencing the LLM’s failure to correctly identify ’no plan’ scenarios varied depending on the model and the prompt style (Tables LABEL:tab:lrnoplanlong and LABEL:tab:lrnoplanshort in Appendix). However, no single variable consistently predicted failure across all models or prompt styles.
Figure 5: Comparison of large language model performance (accuracy) in detecting ’no plan’ scenarios, evaluated with verbose and concise prompt styles
@@ 3.1.4 Feasible Plan
When restricted to problems with feasible plans, Llama-3.1-405B-instruct-fp8 demonstrated the strongest performance, achieving the high rates of generating sound (91%), valid (85%), and optimal plans (69%) using the concise prompt style (Figure 6). I further investigated the influence of problem characteristics – including optimal plan length, number of custom actions, arity, coupling, proportion of slot-fillable variables, and number of goals – on the generation of valid plans. The relationship between these explanatory variables and plan quality (soundness, validity, and optimality) varied considerably across language models.
Figure 6: Performance comparison of large language models on planning tasks with feasible plans
Across most models, optimal plan length was the strongest predictor of valid plan generation (Tables LABEL:tab:lrplanvaliditylong and LABEL:tab:lrplanvalidityshort in Appendix). However, this relationship was not consistent. With the verbose prompt style, logistic regression with L1 regularization ($\alpha=1$) failed to identify any significant predictors of valid plan generation for Granite-3.3-8B-instruct and Llama-3.1-8B-instruct. Likewise, when using the concise prompt, optimal plan length was not a significant predictor for Llama-3.3-70B-instruct, Mixtral-8x22B-instruct-v0.1, and Qwen2.5-72B-instruct. These results indicate that the influence of problem difficulty on valid workflow generation is model-dependent.
Regarding optimal plan generation, coupling consistently explained a significant portion of the variance across all models when the verbose prompt was used (Table LABEL:tab:lrplanoptimallong in Appendix). Meanwhile, this trend was observed only in DeekSeek-V3 and Llama-3.3-70B-instruct when using the concise prompt style (Table LABEL:tab:lrplanoptimalshort in Appendix). These results suggest that the impact of problem characteristics on optimal plan generation is both model- and prompt-dependent.
@@ 3.2 Natural Language to JSON Translation
Llama-3.1-405B-instruct-FP8 achieved the highest success rate in generating valid JSON strings (99%) and minimizing errors in content (Figure 7). Specifically, 73% of the JSON outputs from this model contained all required elements without hallucinated information. Meanwhile, Codellama-34B-instruct-HR and Llama-3.3-70B-instruct consistently failed to generate valid JSON. Across all models, the primary source of translation error was the omission of required elements, rather than the inclusion of hallucinated information. Linear regression analysis revealed a negative correlation between the proportion of correctly translated elements and both the number of custom actions and their arity (Table LABEL:tab:olstranslation in Appendix). Conversely, a positive correlation was observed between correct translation and both the degree of coupling between custom actions and the proportion of slot-fillable variables.
@@
4 Related Work
Current evaluation of reasoning capacity often focuses on final answer accuracy for benchmarks like MATH500 [7], AIME24, GSM8K [8], and MATH. However, this metric overlooks the validity of the reasoning ’process’ itself; a correct answer can be achieved through flawed reasoning, effectively rewarding a successful, yet invalid, plan. Furthermore, relying solely on final answer correctness leaves the quality and structure of the reasoning process opaque. Similarly, BIRD [9], a natural language to SQL translation benchmark, primarily relies on final answer accuracy, despite the inherent planning aspects of SQL query construction alongside elements like intent detection. GSM-Symbolic [10] attempts to address this by generating mathematical problems from symbolic templates, allowing for manual verification of solutions. ReasonAgain [11] uses GPT-4o to generate Python programs for GSM8K and MATH, validating reasoning by executing the generated code.
Beyond these benchmarks, several studies use “puzzle” datasets [12], which include Tower of Hanoi, river crossing, blocks world, and checker jumping, to assess LLM reasoning. These puzzles offer the advantage of verifiable reasoning steps through simulation and can be scaled to challenge LLMs as they saturate. However, the applicability of these abstract puzzles to real-world scenarios remains limited.
A growing body of work focuses on evaluating and training LLMs for workflow automation and reasoning. Several datasets have been proposed to assess LLM planning capabilities, differing in their generation methods (static vs. dynamically generated) and evaluation approaches (manual vs. programmatic). This section details relevant work, highlighting these key distinctions.
4.1 Static Datasets & Primarily Manual/LLM-as-Judge Evaluation
FlowBench [13] presents a static dataset of 51 scenarios across 6 domains, focusing on multi-turn interactions. Evaluation relied heavily on human verification by comparing expected and observed traces, identifying a need for improved LLM planning. Similarly, HuggingGPT [14] uses a static, human-annotated dataset augmented with GPT-4 generated pseudo-labels, and uses LLM-as-a-judge for graph tasks. TaskLAMA [15] also provides a static dataset, created through human annotation, for structured complex task decomposition. ISG-BENCH [16] focuses on interleaved text and image generation, using a static dataset and relying on MLLM-as-a-judge for scoring alignment, coherence, and helpfulness, grounded in human-annotated truth. WikiPlan and RecipePlan contribute static text-image pairs for task completion. These datasets, while valuable, often require significant human effort for creation and/or evaluation [17].
4.2 Static Datasets & Programmatic Evaluation
TaskBench [18] divides task automation into decomposition, tool selection, and parameter prediction, leveraging a static dataset of 17,331 samples generated with GPT-4. The authors used programmatic metrics (F1, Normalized Edit Distance, Rouge) for evaluation, with some partial manual inspection. This approach offers scalability due to a large sample size but can be limited by the quality and coverage of the initial dataset. Open Grounded Planning [2] also aims to generate executable plans, providing LLMs with an extensive action library and leveraging existing datasets. Evaluation combines programmatic metrics (executability, pass rate) with LLM-as-a-judge for assessing completeness, feasibility, and relevance.
4.3 Algorithmically Generated Datasets & Programmatic Evaluation
My work builds upon these efforts by presenting a fully automated pipeline for both data generation and evaluation, addressing the limitations of static datasets and the need for scalable, programmatic (symbolic) assessment of LLM workflow generation capability.